how do you solve the following problems?
24s = 144
(10m)x 21 = 1890
36 = 3k x4
5x = 225/15
3p = 300
(42)n = 48
4 squared n = 48
Please show work.

Answer by CubeyThePenguin(3113)

24s = 144
s = 144/24 = 6

(10m)(21) = 1890
210m = 1890
m = 1890/210 = 9

36 = (3k)(4)
36 = 12k
k = 36/12 = 3

5x = 225/15
5x = 15
x = 15/5 = 3

3p = 300
p = 300/3 = 100

42n = 48
n = 48/42 = 8/7

(4^2)n = 48
16n = 48
n = 48/16 = 3